OpenStudy (anonymous):

Isnt 9*10(to the third power as an exponent)-2000 equal to 7000?

OpenStudy (anonymous):

do i use pemdas?

OpenStudy (anonymous):

=9*10^3 -2000 =9000-2000 =7000